OpenStudy (marie.ella):

Compare 9 ⋅ 10^4 to 3 ⋅ 10^2.

jhonyy9 (jhonyy9):

9 * 10^4 > 3*10^2 hope this will help you

OpenStudy (zzr0ck3r):

Hint: \((3*10^2)^2=3^2(10^2)^2=9*10^4\)

OpenStudy (sshayer):

9*10^4>3*10^4>3*10^2